In a diabetes study on 140 diabetic patients, 25% had uncontrolled blood glucose level. Construct the 95% CI for the population proportion. (a = 0.05)
Answer #1

Answer

Given that the sample proportion is p(hat) = 25/100 = 0.25

sample size n = 140

z critical for 95% confidence interval is 1.96

Confidence interval =

P 2 * ((1-P) *P)/n = 0.25 +1.96 * ((1 -0.25) * 0.25)/140 = 0.25 + 1.96 * 0.0366 = (0.1783, 0.3217)

Required 95% confidence interval for the population proportion is (0.1783,0.3217)